16.

On April 1, the price of gas at Bobs Corner Station was $4.60 per gallon. On May 1, the price was $5.35 per gallon. On June 1, it was back down to $4.60 per gallon. Between April 1 and May 1, Bobs price increased by or Between May 1 and June 1, Bobs price decreased by or Suppose that at a gas station across the street, prices are always 20% higher than Bobs. In absolute dollar terms, the difference between Bobs prices and the prices across the street is when gas costs $5.35 than when gas costs $4.60. Some economists blame high commodity prices (including the price of gas) on interest rates being too low. Suppose the Fed raises the target for the federal funds rate from 2% to 2.75%. This change of raised its target by approximately percentage points means that the Fed
Answer #1

1) Increase in price = Price in May 1 - Price in April 1 = 5.35 - 4.60 = 0.75

or Increase in price = (Price in May 1 - Price in April 1)/Price in April 1 = (5.35 - 4.60)/4.60 x 100 = 0.75/4.60 x 100 = 16.30%

2) Decrease in price = Price in June 1 - Price in May 1 = 4.60 - 5.35 = - 0.75

Decrease in price = (Price in June 1 - Price in May 1)/Price in May 1 = - 0.75/5.35 x 100 = - 14.02%

3) When price at Bob's corner is 4.60 then price in street = 4.60 + 20% of 4.60 = 4.60 + 0.92 = 5.52

When price at Bob's corner is 5.35 then price in street = 5.35 + 20% of 5.35 = 5.35 + 1.07 = 6.42

Difference = 6.42 - 5.35 = $ 1.07

4) (2.75 - 2) = 0.75%